# Leave-One-Out Neighborhood Smoothing for Graphons: Berry-Esseen Bounds, Confidence Intervals, and Honest Tuning

### 28 Feb 2026 v4

@@ 3.4.2 Normal‑approximation intervals
4 Honest tuning via cross-validation
We use the term “honest” to denote that no held-out edge $A_{ij}$ is used in constructing its predictor $\tilde{P}_{ij}^{(-j)}$. This eliminates double use of data but does not imply oracle optimality of the selected bandwidth.
4.1 Cross-validation procedure and notation
$\mathcal{G}\subset\{1,2,\dots,n-2\},$
Define the LOO cross-validation score
Because the neighborhood $\mathcal{N}_{i}^{(-j)}(h)$ is constructed from $A^{(-j)}$ (which omits all edges incident to vertex $j$) and $i\notin\mathcal{N}_{i}^{(-j)}(h)$ by construction, the held‑out edge $A_{ij}$ is conditionally independent of the predictor $\tilde{P}_{ij}^{(-j)}(h)$ given $P$. Consequently, $\widehat{R}_{i}(h)$ is (conditionally) unbiased for the corresponding prediction risk. A full oracle inequality analysis of this tuning rule is not included in this version.
The corresponding prediction risk is the conditional expectation of the same quantity. Because $A_{ij}$ is conditionally independent of $\mathcal{N}_{i}^{(-j)}(h_{n})$ given $P$, this evaluates to
$R_{i}(h_{n}):=\frac{1}{n-1}\sum_{j\neq i}\mathbb{E}\!\left[\bigl(A_{ij}-\tilde{P}_{ij}^{(-j)}(h_{n})\bigr)^{2}\,\middle|\,\mathcal{N}_{i}^{(-j)}(h_{n}),P\right].$ (4.1)
We then select any minimizer of the empirical score over the grid,
$\hat{h}_{n}\in\arg\min_{h_{n}\in\mathcal{G}}\widehat{R}_{i}(h_{n}).$
5 Simulation Study
@@ 5.1 Setup
The empirical coverage levels reported in Table 2 are higher than typical for $n=200$. This reflects the favorable latent geometry and relatively generous bias cushion used in our simulation design. In more challenging settings, we expect coverage to be closer to the nominal level.
For each graphon, we generated an undirected graph with $n=200$ nodes by sampling latent positions $\xi_{i}\sim\mathrm{Unif}[0,1]$ and edges $A_{ij}\sim\mathrm{Bernoulli}(f(\xi_{i},\xi_{j}))$ independently for $i<j$. The neighborhood size and trimming quantile were fixed at
@@
The classical ZLZ smoother constructs neighborhoods using the full adjacency matrix. The LOO smoother recomputes neighborhoods using $A^{(-j)}$, the adjacency matrix with column and row $j$ removed, thereby restoring conditional independence between the neighborhood and the target edge. Both estimators were symmetrized via $\widehat{P}_{ij}=(\tilde{P}_{ij}+\tilde{P}_{ji})/2$.
@@ 5.2 Evaluation metrics
    $w_{ij}^{(B)}=\sqrt{2\widehat{V}_{ij}\log(4/\alpha)}+\frac{\log(4/\alpha)}{3h},$
@@ 5.3 Results
Graphon $\mathrm{MSE}_{0}^{\mathrm{LOO}}$ $\mathrm{MSE}_{0}^{\mathrm{ZLZ}}$ $\mathrm{Cov}_{\mathrm{Bousquet}}$ $\mathrm{Cov}_{\mathrm{Normal}}$
Smooth 0.0027 0.0030 0.999 0.995
Block 0.0021 0.0027 1.000 1.000
Wiggle 0.0031 0.0030 0.944 0.885
Rank-one 0.0029 0.0030 0.999 0.996
Spiky 0.0021 0.0033 0.977 0.977
Table 2: Row-wise MSE and empirical coverage for five graphons at $n=200$.
@@ 6 Discussion and Extensions
    Extension to sparse regimes. For models of the form $P_{ij}=\rho_{n}W(\xi_{i},\xi_{j})$ with $\rho_{n}\to 0$, the effective variance of $U_{ij}^{(-j)}$ shrinks with $\rho_{n}$. Normal approximation would require renormalization by $\sqrt{h_{n}\rho_{n}}$, and empirical Bernstein bounds must account for rare-event behavior. While the leave-one-out decoupling remains valid, the concentration and Berry-Esseen analysis would require substantially different arguments.

### 14 Mar 2026 v5

@@ 6 Simulation Study
To comprehensively evaluate the proposed uncertainty quantification, we track three primary metrics across the varying graphon topologies: the mean squared error (MSE) for point estimation, the empirical coverage rate, and the average confidence interval width. While empirical coverage verifies the theoretical validity of our bounds, tracking the average interval width explicitly quantifies the inferential cost. Specifically, comparing the average bounds allows us to explicitly demonstrate the trade-off between the conservative, strictly non-asymptotic Empirical Bernstein intervals and the sharper, asymptotically valid Normal approximations. The results of this numerical study are summarized in Table 2.
@@ 6.1 Setup
The empirical coverage levels reported in Table 2 are higher than typical for $n=500$. This reflects the favorable latent geometry and relatively generous bias cushion used in our simulation design. In more challenging settings, we expect coverage to be closer to the nominal level.
For each graphon, we generated an undirected graph with $n=500$ nodes by sampling latent positions $\xi_{i}\sim\mathrm{Unif}[0,1]$ and edges $A_{ij}\sim\mathrm{Bernoulli}(f(\xi_{i},\xi_{j}))$ independently for $i<j$. The neighborhood size and trimming quantile were fixed at
@@
The classical ZLZ smoother constructs neighborhoods using the full adjacency matrix. The LOO smoother recomputes neighborhoods using $A^{(-j)}$, the adjacency matrix with column and row $j$ removed, thereby restoring conditional independence between the neighborhood and the target edge. To evaluate point estimation accuracy, both estimators were symmetrized via $\widehat{P}_{ij}=(\tilde{P}_{ij}+\tilde{P}_{ji})/2$ for the MSE calculations. However, to strictly adhere to our theoretical independence guarantees, the confidence intervals were constructed and evaluated using the fundamental one-sided predictor $\tilde{P}_{ij}$.
@@ 6.2 Evaluation metrics
    $w_{ij}^{(B)}=\sqrt{\frac{2s_{ij}^{2}\log(4/\alpha)}{h_{n}}}+\frac{7\log(4/\alpha)}{3(h_{n}-1)}.$
@@ 6.3 Results
Table 2: Finite-sample performance of the LOO estimator across various graphon topologies ($n=500$). We report the mean squared error (MSE) for both the asymmetric LOO and classical symmetrized estimators for a fixed row. Coverage and average interval widths are compared between the strictly non-asymptotic Empirical Bernstein bounds and the asymptotic Normal approximation.
Point Estimation (MSE) Uncertainty Quantification (LOO)
Graphon LOO Classical Cov. (EB) Width (EB) Cov. (Normal) Width (Normal)
Smooth $0.0019$ $0.0020$ $1.000$ $0.554$ $0.980$ $0.272$
Block $0.0014$ $0.0020$ $1.000$ $0.539$ $0.985$ $0.263$
Wiggle $0.0061$ $0.0060$ $0.999$ $0.557$ $0.919$ $0.272$
Rank-1 $0.0015$ $0.0016$ $1.000$ $0.530$ $0.983$ $0.264$
Spiky $0.0005$ $0.0009$ $1.000$ $0.440$ $0.989$ $0.230$
The results in Table 2 provide strong empirical validation for both the estimation efficiency and the inferential validity of the leave-one-out framework.
First, regarding point estimation, the asymmetric LOO predictor $\tilde{P}_{ij}$ demonstrates highly competitive mean squared error compared to the classical symmetrized estimator $\widehat{P}_{ij}$. In fact, the LOO estimator matches or slightly improves upon the classical MSE in four of the five tested topologies. This confirms our theoretical assertion that strictly excluding the target column during neighborhood formation incurs a negligible finite-sample efficiency penalty, preserving the minimax optimal rate while unlocking conditional independence.
Second, the coverage results clearly illustrate the theoretical trade-offs between our two inferential approaches. The Empirical Bernstein intervals, grounded in strictly non-asymptotic bounds, achieve absolute coverage (near $1.000$) across all topologies. However, this theoretical guarantee comes at the cost of practical sharpness, yielding intervals that span the majority of the $[0,1]$ support. In contrast, the asymptotic intervals constructed via our Central Limit Theorem cut the average interval widths by more than half (averaging roughly $0.25$). Crucially, these Normal approximations successfully maintain the nominal $0.95$ coverage rate on the smooth, block, rank-1, and spiky graphons at sample size of $n=500$.
Finally, the wiggle graphon provides an instructive boundary case regarding finite-sample deterministic bias. While the Normal approximation achieves $0.919$ coverage, the under-coverage relative to the $0.95$ nominal target is a direct consequence of the graphon’s high-frequency gradients ($\sin(4\pi uv)$). Such highly variable local geometries inflate the deterministic approximation error at finite $n$.
@@ 7 Discussion and Extensions
    Extension to sparse regimes. For models of the form $P_{ij}=\rho_{n}f(\xi_{i},\xi_{j})$ with $\rho_{n}\to 0$, the effective variance of $U_{ij}^{(-j)}$ shrinks with $\rho_{n}$. Normal approximation would require renormalization by $\sqrt{h_{n}\rho_{n}}$, and empirical Bernstein bounds must account for rare-event behavior. While the leave-one-out decoupling remains valid, the concentration and Berry-Esseen analysis would require substantially different arguments.